OpenStudy (anonymous):

Can I know how to solve this problem????The question is 4^x/16=64/2^y

OpenStudy (akashdeepdeb):

Hey @Babypanda What do you have to solve for in this question? :)

OpenStudy (anonymous):

WELL MY ANSWER WAS x=3 AND Y=4 but my friend said the answer is WRONG!!!!

OpenStudy (anonymous):

AkashdeepDeb are you having a deep thought about the question??????

OpenStudy (akashdeepdeb):

Try putting those values into the equation, if you get the answer, you are right, if you don't, your friend is. Go ahead, like this: \[\frac{4^x}{16} = \frac{64}{2^y}\] Cross multiply. \[4^x . 2^y = 64.16\] NOTE: [The dot implies multiplication] Now express this in powers of '2'. Because all the numbers you see are powers of 2! :D \[2^{2x}~.~2^y = 2^6~.~2^4\] Now let's use some exponents' rules. We know that, \(~~~a^x~.~a^y = a^{x+y}\) So, \(~~~a^{any~number}~.~a^{some ~number} = a^{any~ number~ + ~some ~number}\) Thus, then our equation becomes, \[2^{2x+y} = 2^{10}\] Now, since the base [2] is the same, even the exponent MUST be the same, right? Becuase, if \(~~~a^(x) = a^(y)\) then \(x\) must be equal to \(y\). :D So, we get, 2x + y = 10. Now if we have another equation for x and y, we can find the value of x,y. Understood this? :)

OpenStudy (akashdeepdeb):

By the way - That is, \[a^x = a^y,~ ~then~ ~x~~ must ~~be~~ equal~~ to~~ y\]

OpenStudy (anonymous):

Thank you so much!!!!!! The question that links to it is log x ( y+2)=1 + log x 4

OpenStudy (akashdeepdeb):

Well, you can another equation from that, do you know how to solve that? ^

OpenStudy (akashdeepdeb):

...can (get) another...

OpenStudy (anonymous):

Is it y=4x -2....The answer for the second equation...

OpenStudy (akashdeepdeb):

Is this the equation: \(~~~log~{x(y+2)} = 1 + log~(x^4)\) ?

OpenStudy (anonymous):

NO!!!!It is log x\[\log_{x}(y+2)=1+\log_{x}4 \]

OpenStudy (akashdeepdeb):

One sec...

OpenStudy (anonymous):

OKAY

OpenStudy (akashdeepdeb):

y = 4x - 2 is right. :) Now solve the equations and see what you get.

OpenStudy (anonymous):

Is it x = 6 and y= 22????

OpenStudy (akashdeepdeb):

No. Check again. That does not satisfy the first equation. :) x=2 y=6

OpenStudy (anonymous):

Yeah !!!! YOU ARE RIGHT !!!! i WROTE THE WRONG EQUATION..... THANK YOU!!!!

OpenStudy (anonymous):

ARE YOU A FORM 5 STUDENT????

OpenStudy (akashdeepdeb):

I am not sure what that is. Yes, but I am a student.

OpenStudy (anonymous):

WAIT , you are not from Malaysia???

OpenStudy (akashdeepdeb):

Haha, no.

OpenStudy (anonymous):

So,all this while I was talking to a foreigner...Cool..

OpenStudy (anonymous):

Well,in our country form 5 means you are 17 years old teenager.....

OpenStudy (akashdeepdeb):

Informative. Thanks for letting me know. I am 16 years old.

OpenStudy (anonymous):

Well, you as same age as me... Nice to know you genius